Find Y' given x + xy + y = 2
1
Expert's answer
2019-09-27T13:52:21-0400

"x+xy+y=2\\
y={\frac {(2-x)} {(x+1)}}\\
y'=({\frac {(2-x)} {(x+1)}})'={\frac {(2-x)'*(x+1)-(2-x)*(x+1)'}{(x+1)^2}}=\\
={\frac {(-1)*(x+1)-(2-x)*1}{(x+1)^2}}={\frac{-3}{(x+1)^2}}"

answer: "{\frac{-3}{(x+1)^2}}"